OpenStudy (anonymous):

the derivative function of y = ax^2 + bx + c is y = 4x + 2. the function has a zero at x = 1. find the values of a, b and c

OpenStudy (anonymous):

a=2 b=2 c=1? I think.

OpenStudy (anonymous):

first find the derivative of ax^2 + bx + c and compare with y' = 4x + 2 : y'=2ax+b then we know that y' = 4x + 2 therefore 2a=4 so a=2 , then b = 2 Also they say that y=0 at x=1 which means: f(x)=y=ax^2 + bx + c f(1)= a(1)^2 + b(1) + c 0= a + b + c plug in the values of a=2 and b=2 in the function 0= 2 + 2 + c c=-4 therefore the values are: a=2 b=2 c=-4